What is the smallest number that has a remainder of 0 when divided by 16 and 36? I have listed the multiples of these two numbers out to see if they have a LCM, but I cannot seem to find one. I know that you can multiply if the two numbers are prime numbers to solve, but in this case they are not. Please help! I need this by tomorrow!
Answer by jim_thompson5910(35256)

GCF of 16 and 36 is 4

LCM = 16*36/(GCF) = 576/4 = 144

So the LCM of 16 and 36 is 144